OpenStudy (anonymous):

Solve the system of equations by substitution. 9x+7y=31, x=54-8y

OpenStudy (anonymous):

replace the x in 9x with (54-8y) So the equation would now read 9(54-8y)+7y=31

OpenStudy (anonymous):

Then isolate the y to one side and you will have your answer as expressed in terms of y

OpenStudy (anonymous):

9(54-8y)+7y=31 486-72y+7y=31 486-65y=31 solve for y

OpenStudy (anonymous):

do you divide 65 by both sides?

OpenStudy (anonymous):

not yet. subtract 486 in both sides first 486-486-65y=31-486

OpenStudy (anonymous):

which would be 455-65y=31 right? then you subtract 65?

OpenStudy (anonymous):

if you subtract a positive 486 from a negative 486 is 0, 31-486= -455 -65y=-455 then you have to solve for y by dividing 65 in both sides

OpenStudy (anonymous):

so then y=7?

OpenStudy (anonymous):

you got it! good job!

OpenStudy (anonymous):

then if you want to solve for y you plug 7 where the y is in the original equation and do the same

OpenStudy (anonymous):

would it be 7=54-8y?

OpenStudy (anonymous):

no, you put it where the y is. x=54-8(7)

OpenStudy (anonymous):

because you are solving for x

OpenStudy (anonymous):

would the whole answer be (-2,7)?

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

Thank you for your help :)

OpenStudy (anonymous):

no problem!